OpenStudy (anonymous):

What is the solution to the equation 49^x = sqrt7^ (2x + 6)

OpenStudy (anonymous):

\[49^{x}=\sqrt{7}^{2x+6}\]

jhonyy9 (jhonyy9):

49^x =7^2x can you rewrite the right part in the same forme ?

OpenStudy (anonymous):

x=3

jhonyy9 (jhonyy9):

so like 7 on what exponent ? and after this you need just make these exponents equales and from this you need calcule the x

jhonyy9 (jhonyy9):

sqrt7 = 7^(1/2) can you continue ?